OpenStudy (anonymous):

Find the exact value of the radical expression in simplest form. the square root of s to the eighth power + the square root of 25 times s to the eighth power + 2 the square root of s to the eighth power − the square root of s to the fourth power 8s4 28 times s to the fourth power times the square root of s 8s4 − s2 28 times s to the fourth power times the square root of s − s2

jhonyy9 (jhonyy9):

can you rewrite it numericaly

jhonyy9 (jhonyy9):

?

OpenStudy (anonymous):

sqrt s^8 + sqrt 25s^8 + 2 sqrt s^8 - sqrt s^4

jhonyy9 (jhonyy9):

ok step by step sqrt s^8 = sqrt (s^4)^2 = ?

jhonyy9 (jhonyy9):

and you know sqrt25 = ?

OpenStudy (cggurumanjunath):